at a certain university 540 students are math and computer science majors. if 300 are majoring in math and 120 are majoring in both, how many are majoring in computer science?
Answer by Edwin McCravy(20086)

Let x = The number of students majoring in math ONLY

let y = The number of students majoring in computer science ONLY

Let z - The number of students majoring in computer science.

 +  = 


 +  =


 +  + = 

So we substitute the corresponding numbers or unknowns for
the above three statements and we have this system of equations:

x + 120 = 300
y + 120 = z
x + 120 + y = 540

Solve that and get x = 180, y = 240, and z = 360

The answer asked for was z, which is 360.

Edwin